OpenStudy (christos):

Calc 2, Can you please tell me why is this considered a 'geometric Serie" instead of arithmetic ?

OpenStudy (christos):

@AllTehMaffs

OpenStudy (turingtest):

could you post the series please?

OpenStudy (christos):

It's number 7 , I am sorry that I forgot to mention this .

OpenStudy (turingtest):

number 7 ? I'm afraid that nobody here is reading your particluar book at the moment, so you'll have to actually write out the series ;)

OpenStudy (christos):

here its this eerie

OpenStudy (christos):

serie*

OpenStudy (kc_kennylau):

this is a series?

OpenStudy (kc_kennylau):

you can break \(\dfrac1{(k+2)(k+3)}\) into \(\dfrac1{k+3}-\dfrac1{k+2}\) and expand the summation

OpenStudy (kc_kennylau):

you'll find that the terms cancel out each other

OpenStudy (christos):

isn't this a sum/sequence ? Or am I wrong ? It's my first day in this chapter so parton me if I make some horrible mistakes , I wish its ok :/

OpenStudy (kc_kennylau):

oh wait sorry my fault this is a series

OpenStudy (christos):

I already did that , but what is next

OpenStudy (christos):

after expanding it *

OpenStudy (kc_kennylau):

and i would say that this is neither an arithmetic series nor a geometric series

OpenStudy (kc_kennylau):

(P.S. and the singular form of series is, you guessed it, still series)

OpenStudy (christos):

and why are you saying its neither of those? then what is it ? can you please tell me how did you find out this?

OpenStudy (kc_kennylau):

so the terms will cancel out each other and you'll be left with \(\dfrac1{\infty+2}-\dfrac13\) which is actually just \(\dfrac13\)

OpenStudy (kc_kennylau):

an arithmetic sequence has to have a common difference while a geometric sequence has to have a common ratio

OpenStudy (kc_kennylau):

i mean \(-\dfrac13\) sorry

OpenStudy (kc_kennylau):

do you need the steps or you can actually do it myself or you want me to guide you instead of just giving you all the steps

OpenStudy (christos):

I prefer to guide me of its possible bro :D

OpenStudy (christos):

But I am really confused

OpenStudy (christos):

I can understand that a sequence can have a same difference between its term

OpenStudy (turingtest):

the series sums to 1/2, but I thought the question was just why it's geometric, i.e. what's the common ratio?

OpenStudy (christos):

but what do you mean by different ratio ? basically what exactly is ratio ?

OpenStudy (kc_kennylau):

lemme write down what we have now, we have: \[\Large\sum_{k=1}^\infty\frac1{(k+2)(k+3)}=\sum_{k=1}^\infty\left(\dfrac1{k+3}-\dfrac1{k+2}\right)\]

OpenStudy (christos):

I have the solution and it says that it sums to ⅓ , but I don't know how tha process took place

OpenStudy (kc_kennylau):

ratio of a and b is \(\dfrac ba\)

OpenStudy (christos):

aaah

OpenStudy (christos):

ok so far I have come across, I even expanded the s eirie ,but after that , I have to admit that I am completely clueless

OpenStudy (kc_kennylau):

sorry i made a mistake there lemme write it again

OpenStudy (zarkon):

\[\frac{1}{(k+2)(k+3)}=\frac{1}{k+2}-\frac{1}{k+3}\]

OpenStudy (kc_kennylau):

we have: \[\large\sum_{k=1}^\infty\frac1{(k+2)(k+3)}=\sum_{k=1}^\infty\left(\frac1{k+2}-\frac1{k+3}\right)\]

OpenStudy (christos):

yes

OpenStudy (kc_kennylau):

and the answer should be \(\dfrac13\) instead of 0.5

OpenStudy (christos):

yes but how, the processs s what I really need to know and be able to reproduce

OpenStudy (turingtest):

sorry, it is 1/3

OpenStudy (anonymous):

This is called a telescoping series and the methods described above are the right way to do it. The answer is 1/3

OpenStudy (kc_kennylau):

expand that and we have: \[\frac1{1+2}-\frac1{1+3}+\frac1{1+3}-\frac1{1+4}+\cdots+\frac1{\infty+2}-\frac1{\infty+3}\]

OpenStudy (kc_kennylau):

cancel them and we got \(\dfrac1{1+2}\) which is \(\dfrac13\)

OpenStudy (christos):

what's the "k" here ? http://screencast.com/t/ojwnBozWwZ

OpenStudy (kc_kennylau):

1

OpenStudy (anonymous):

The formal way of doing it is to compute the partial sum \[ s_n=\sum _{k=1}^n \frac{1}{(k+2) (k+3)}=\frac{1}{3}-\frac{1}{n+3} \] And let n go to infinity

OpenStudy (christos):

How did you find out this part in one step http://screencast.com/t/6K4sooTrfU5 @eliassaab

OpenStudy (anonymous):

I cheated, all the remaining terms cancel as described above/

OpenStudy (kc_kennylau):

To determine whether it's an arithmetic sequence or a geometric sequence, we consider the first three terms: Difference of first two terms\(=\dfrac1{3\times4}-\dfrac1{4\times5}=\dfrac1{30}\) Difference of second two terms\(=\dfrac1{4\times5}-\dfrac1{5\times6}=\dfrac1{60}\) Since they are not equal, it is not an arithmetic sequence. Ratio of the first two terms\(=\dfrac1{4\times5}\div\dfrac1{3\times4}=\dfrac35\) Ratio of the second two terms\(=\dfrac1{5\times6}\div\dfrac1{4\times5}=\dfrac23\) Since they are not equal, it is not a geometric sequence.

OpenStudy (anonymous):

I do not want to repeat what the others have helped with so far.

OpenStudy (christos):

so if its not a geographic neither an arithmetic the its a telescopic ?

OpenStudy (christos):

then*

OpenStudy (zarkon):

no

OpenStudy (kc_kennylau):

it can be neither of the three

OpenStudy (christos):

ok I wil practice this and I will learned the technique, thanks so much kc_kennylau for your patience and help

OpenStudy (kc_kennylau):

no problem :) give credits to others too :)

OpenStudy (christos):

ofc !, thank you so much @TuringTest @eliassaab @Zarkon

OpenStudy (anonymous):

Here all the steps to compute the partial sum \[ s_n=\sum _{k=1}^n \frac{1}{(k+2) (k+3)}=\sum _{k=1}^n \left(\frac{1}{k+2}-\frac{1}{k+3}\right)=\\\frac{1}{3}+\frac{1}{4}-\frac{1}{4}-\frac{1}{5}+ \cdots -\frac{1}{n+2}+\frac{1}{n+2}-\frac{1}{n+3}=\frac{1}{3}-\frac{1}{n+3} \]

OpenStudy (anonymous):

YW